If 4 is added to 7 times a positive number, the result is twice the square of the number. Find the number.
Answer by ewatrrr(24785)

7n + 4 = 2n^2
2n^2 - 7n - 4 = 0
(2n +1)(n - 4)= 0
n = 4 (a positive number)
And...
28 + 4 = 2*16 CHECKS!